OpenStudy (anonymous):

find horizontal and vertical asymptotes of f(x)= 2x^2+6x-20 / 3x^2 -36x+60

OpenStudy (anonymous):

horizontal asymptote y=2/3

OpenStudy (anonymous):

thank you and how do i get the vertical asymptote

OpenStudy (anonymous):

\[\frac{2 x^2+6 x-20}{3 x^2-36 x+60}=\frac{2 (x+5)}{3 (x-10)} \]Refer to the plot attachment.